OpenStudy (anguyennn):

How go from a contract logarithm form to expanded logarithm form? log(3x-3y)

zepdrix (zepdrix):

Hey there :) \[\rm \log[3x-3y]\]

zepdrix (zepdrix):

Let's start by factoring a 3 out of each the x and the y,\[\rm \log[3(x-y)]\]

OpenStudy (anguyennn):

yes

OpenStudy (anguyennn):

Thank you for helping me btw!

OpenStudy (anguyennn):

very much appreciated

zepdrix (zepdrix):

Recall your Log Product Rule:\[\rm \log(a\cdot b)=\log(a)+\log(b)\]

zepdrix (zepdrix):

Do you see how we can apply that to our problem maybe?\[\rm \log[3\cdot(x-y)]\]Do you see which two things are being multiplied together?

OpenStudy (anguyennn):

yes the 3 and (x-y)

zepdrix (zepdrix):

Good :) So how is that going to break down using the rule? Can you figure it out?

OpenStudy (anguyennn):

log(3)+log(x-y)

zepdrix (zepdrix):

Ah yes! Good job! That's pretty much all we can do with this one.

OpenStudy (anguyennn):

oh!!! really!!!

OpenStudy (anguyennn):

thats it?

zepdrix (zepdrix):

The x and the y are being subtracted, we don't have a special rule for that unfortunately.

OpenStudy (anguyennn):

there were a couple different ones i was stuck on as well

zepdrix (zepdrix):

k let's check em out :)

OpenStudy (anguyennn):